OpenStudy (anonymous):

If ⅜ of a 60-pound bag of ready-mix concrete is Portland cement, how many pounds of other materials are in the bag?

OpenStudy (texaschic101):

3/8 of 60 lbs is Portland cement... 3/8 * 60 = 180/8 = 22 1/2 lbs. other materials.. 60 - 22 1/2 = 60 - 22.5 = 37.5 (or 37 1/2) 22 1/2 lbs is Portland cement and 37 1/2 lbs is other material.